Name:

Philosophy 29—Critical Reasoning Midterm Exam A November 5, 1997

True/False (2 pts. each)

Print “T” if the statement is true, otherwise print “F” A proposition and its double negation are logically equivalent. For purposes of evaluating the credibility of a source, a statement is less credible if the statement

maker knows that his or her reputation is at risk.

Observation reports by the observer are generally to be preferred over others’ reports of these observation reports.

Observations and the conclusions inferred from them are usually equally reliable. A conclusion can also be a reason for another conclusion. A deductively valid argument can have true premises and a true conclusion. If an argument is bad, the conclusion must be wrong. A deductively valid argument can have false premises and a true conclusion. The word “because” usually indicates that the next sentence is the conclusion of an argument.

Section 2. Fill in the blanks (3 pts. each)

1. The then part of a conditional is called the .

2. When examining an argument, there are three things you must do. These are:

(a) (b) (c)

3. Write the contrapositive of the proposition: If John is tall, then John is qualified for the basketball team.

4. An argument is deductively invalid if

5. What fallacy labels best describe the following?

(a) You are giving reasons why the referee’s decision was a bad one because you are for the Lakers. So I don’t accept your conclusion.

(b) Carl Sagan says, “Asics are the best shoes to wear when thinking about astronomy.”

6. Write whether each of the following is an observation or a conclusion:

(a) The physicist says, “This is the particle track of an electron.”

(b) The physicist says, “The particle track on this film curves to the right.”

7. If an argument has a conditional proposition as a premise and the affirmation of the antecedent of the conditional as a premise, then it is deductively valid to conclude

Answer Key for Exam A

True/False (2 pts. each)

Print “T” if the statement is true, otherwise print “F” \overline{True} A proposition and its double negation are logically equivalent. \overline{True} For purposes of evaluating the credibility of a source, a statement is less credible if the statement

maker knows that his or her reputation is at risk. \overline{True} Observation reports by the observer are generally to be preferred over others’ reports of these

observation reports. \overline{False} Observations and the conclusions inferred from them are usually equally reliable. \overline{True} A conclusion can also be a reason for another conclusion. \overline{True} A deductively valid argument can have true premises and a true conclusion. \overline{False} If an argument is bad, the conclusion must be wrong. \overline{True} A deductively valid argument can have false premises and a true conclusion. \overline{False} The word “because” usually indicates that the next sentence is the conclusion of an argument.

Section 2. Fill in the blanks (3 pts. each)

1. The then part of a conditional is called the .

2. When examining an argument, there are three things you must do. These are:

(a) (b) (c)

3. Write the contrapositive of the proposition: If John is tall, then John is qualified for the basketball team.

4. An argument is deductively invalid if

5. What fallacy labels best describe the following?

(a) You are giving reasons why the referee’s decision was a bad one because you are for the Lakers. So I don’t accept your conclusion.

(b) Carl Sagan says, “Asics are the best shoes to wear when thinking about astronomy.”

6. Write whether each of the following is an observation or a conclusion:

(a) The physicist says, “This is the particle track of an electron.”

(b) The physicist says, “The particle track on this film curves to the right.”

7. If an argument has a conditional proposition as a premise and the affirmation of the antecedent of the conditional as a premise, then it is deductively valid to conclude Name:

Philosophy 29—Critical Reasoning Midterm Exam B November 5, 1997

True/False (2 pts. each)

Print “T” if the statement is true, otherwise print “F”

Observation reports by the observer are generally to be preferred over others’ reports of these observation reports.

Observations and the conclusions inferred from them are usually equally reliable. A deductively valid argument can have true premises and a true conclusion. A conclusion can also be a reason for another conclusion. A proposition and its double negation are logically equivalent. If an argument is bad, the conclusion must be wrong. For purposes of evaluating the credibility of a source, a statement is less credible if the statement

maker knows that his or her reputation is at risk.

The word “because” usually indicates that the next sentence is the conclusion of an argument. A deductively valid argument can have false premises and a true conclusion.

Section 2. Fill in the blanks (3 pts. each)

1. Write the contrapositive of the proposition: If John is tall, then John is qualified for the basketball team.

2. An argument is deductively invalid if

3. The then part of a conditional is called the .

4. If an argument has a conditional proposition as a premise and the affirmation of the antecedent of the conditional as a premise, then it is deductively valid to conclude

5. When examining an argument, there are three things you must do. These are:

(a) (b) (c)

6. What fallacy labels best describe the following?

(a) You are giving reasons why the referee’s decision was a bad one because you are for the Lakers. So I don’t accept your conclusion.

(b) Albert Einstein says, “Reeboks are the best shoes to wear when thinking about physics.”

7. Write whether each of the following is an observation or a conclusion:

(a) The physicist says, “This is the particle track of an electron.”

(b) The physicist says, “The particle track on this film curves to the right.”

Answer Key for Exam B

True/False (2 pts. each)

Print “T” if the statement is true, otherwise print “F” \overline{True} Observation reports by the observer are generally to be preferred over others’ reports of these

observation reports. \overline{False} Observations and the conclusions inferred from them are usually equally reliable. \overline{True} A deductively valid argument can have true premises and a true conclusion. \overline{True} A conclusion can also be a reason for another conclusion. \overline{True} A proposition and its double negation are logically equivalent. \overline{False} If an argument is bad, the conclusion must be wrong. \overline{True} For purposes of evaluating the credibility of a source, a statement is less credible if the statement

maker knows that his or her reputation is at risk. \overline{False} The word “because” usually indicates that the next sentence is the conclusion of an argument. \overline{True} A deductively valid argument can have false premises and a true conclusion.

Section 2. Fill in the blanks (3 pts. each)

1. Write the contrapositive of the proposition: If John is tall, then John is qualified for the basketball team.

2. An argument is deductively invalid if

3. The then part of a conditional is called the .

4. If an argument has a conditional proposition as a premise and the affirmation of the antecedent of the conditional as a premise, then it is deductively valid to conclude

5. When examining an argument, there are three things you must do. These are:

(a) (b) (c)

6. What fallacy labels best describe the following?

(a) You are giving reasons why the referee’s decision was a bad one because you are for the Lakers. So I don’t accept your conclusion.

(b) Albert Einstein says, “Reeboks are the best shoes to wear when thinking about physics.”

7. Write whether each of the following is an observation or a conclusion:

(a) The physicist says, “This is the particle track of an electron.”

(b) The physicist says, “The particle track on this film curves to the right.”

Name:

Philosophy 29—Critical Reasoning Midterm Exam C November 5, 1997

True/False (2 pts. each)

Print “T” if the statement is true, otherwise print “F” The word “because” usually indicates that the next sentence is the conclusion of an argument. For purposes of evaluating the credibility of a source, a statement is less credible if the statement

maker knows that his or her reputation is at risk.

A proposition and its double negation are logically equivalent. A deductively valid argument can have false premises and a true conclusion. Observation reports by the observer are generally to be preferred over others’ reports of these

observation reports.

A conclusion can also be a reason for another conclusion. If an argument is bad, the conclusion must be wrong. A deductively valid argument can have true premises and a true conclusion. Observations and the conclusions inferred from them are usually equally reliable.

Section 2. Fill in the blanks (3 pts. each)

1. If an argument has a conditional proposition as a premise and the affirmation of the antecedent of the conditional as a premise, then it is deductively valid to conclude

2. Write the contrapositive of the proposition: If John is tall, then John is qualified for the basketball team.

3. An argument is deductively invalid if

4. When examining an argument, there are three things you must do. These are:

(a) (b) (c)

5. Write whether each of the following is an observation or a conclusion:

(a) The physicist says, “This is the particle track of an electron.”

(b) The physicist says, “The particle track on this film curves to the right.”

6. The then part of a conditional is called the .

7. What fallacy labels best describe the following?

(a) You are giving reasons why the referee’s decision was a bad one because you are for the Lakers. So I don’t accept your conclusion.

(b) Carl Sagan says, “Asics are the best shoes to wear when thinking about astronomy.”

Answer Key for Exam C

True/False (2 pts. each)

Print “T” if the statement is true, otherwise print “F” \overline{False} The word “because” usually indicates that the next sentence is the conclusion of an argument. \overline{True} For purposes of evaluating the credibility of a source, a statement is less credible if the statement

maker knows that his or her reputation is at risk. \overline{True} A proposition and its double negation are logically equivalent. \overline{True} A deductively valid argument can have false premises and a true conclusion. \overline{True} Observation reports by the observer are generally to be preferred over others’ reports of these

observation reports. \overline{True} A conclusion can also be a reason for another conclusion. \overline{False} If an argument is bad, the conclusion must be wrong. \overline{True} A deductively valid argument can have true premises and a true conclusion. \overline{False} Observations and the conclusions inferred from them are usually equally reliable.

Section 2. Fill in the blanks (3 pts. each)

1. If an argument has a conditional proposition as a premise and the affirmation of the antecedent of the conditional as a premise, then it is deductively valid to conclude

2. Write the contrapositive of the proposition: If John is tall, then John is qualified for the basketball team.

3. An argument is deductively invalid if

4. When examining an argument, there are three things you must do. These are:

(a) (b) (c)

5. Write whether each of the following is an observation or a conclusion:

(a) The physicist says, “This is the particle track of an electron.”

(b) The physicist says, “The particle track on this film curves to the right.”

6. The then part of a conditional is called the .

7. What fallacy labels best describe the following?

(a) You are giving reasons why the referee’s decision was a bad one because you are for the Lakers. So I don’t accept your conclusion.

(b) Carl Sagan says, “Asics are the best shoes to wear when thinking about astronomy.”